Main INDEX
Monthly INDEX
PREV
NEXT
User name lkaufman/paschke
Log entry time 16:19:47 on July18,2004
Entry number 128875
keyword=Hall C IA picomotor adjusted
We moved the S1 picomotor to boost the strength of the Hall C IA. This was necessary to reliably zero the Hall C Aq with IHWP=OUT.
The picomotor was moved from 1000 to 1200.
The new Hall C IA slope is -2100 ppm/Volt. The flipper files for IN and OUT states have been updated with new guesses for the C-IA settle point, and the feedback .db file has been updated with the new slope.